OpenStudy (anonymous):

Find an Euler differential equation of order two with solution: y=x^3

OpenStudy (anonymous):

Eulers sub is t=ln x am i right?

OpenStudy (anonymous):

e^t=x so t=lnx so yep

OpenStudy (anonymous):

well then e^(3t) is an answer for eulers rearranged equation for t

OpenStudy (anonymous):

x^3=(e^t)^3=e^(3t)

OpenStudy (anonymous):

am i clear? i think im not :D

OpenStudy (anonymous):

Yep, perfectly clear

OpenStudy (anonymous):

well for example i choose (m-3)(m+3) for equation (t equation)

OpenStudy (anonymous):

i mean like: \[\large y_{t}''-9y_{t}=0\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

now u just need to convert it to Euler equation (in terms of x)

OpenStudy (anonymous):

and u should know an Euler equation \[x^2 y'' +axy'+by=0\] with t=ln x becomes to \[y''+(a-1)y'+by=0\]

OpenStudy (anonymous):

Ok, yes, right until now

OpenStudy (anonymous):

now u tell me what is the values for a and b

OpenStudy (anonymous):

a=1 b=-9 Am I right?

OpenStudy (anonymous):

yep and what about euler equation?

OpenStudy (anonymous):

x^2y''+xy'-9y=0 xdd

OpenStudy (anonymous):

almost done

OpenStudy (anonymous):

u can get many euler equations by choosing (m-3)(m+k)

OpenStudy (anonymous):

Ok, thanks @mukushla, great as always

OpenStudy (anonymous):

welcome my friend